# TimeBridge: Non-Stationarity Matters for Long-term Time Series Forecasting

### v2

@@ E.2 Ablation Studies
We present the full results of the ablation studies discussed in the main text. Tab. 13 provides the complete results of the ablation on removing non-stationarity in both Integrated and Cointegrated Attention. Tab. 14 reports the full results on the impact and order of Integrated and Cointegrated Attention, with an illustrative visualization in Fig. 6. Additionally, Tab. 15 shows the results of ablation on different modeling approaches for these attention mechanisms. Finally, Tab. 16 presents the results of varying the number of downsampled patches $M$ and its effect on forecasting performance.
@@
$\checkmark$ $\times$ 96 0.143 0.192 0.161 0.224 0.118 0.218 0.340 0.240
192 0.185 0.235 0.177 0.237 0.142 0.237 0.343 0.250
$\checkmark$ $\times$ 336 0.237 0.277 0.188 0.244 0.156 0.252 0.363 0.257
720 0.307 0.330 0.197 0.252 0.179 0.278 0.393 0.271
Avg. 0.218 0.259 0.181 0.239 0.149 0.246 0.360 0.255
@@ Appendix F Statistical Analysis
Appendix G Visualization
Fig. 7 visualizes short-term fluctuations and long-term cointegration across stock sectors. Fig. 8 provides additional examples of intra-variate attention maps comparing stationary and non-stationary conditions for different patches in the Electricity dataset. Fig. 9 shows further examples of inter-variate attention maps in the Solar dataset under both stationary and non-stationary conditions. Fig. 10, Fig. 11, Fig. 12, and LABEL:fig:app_vis_traffic present long-term forecasting visualizations for Weather, Solar, Electricity, and Traffic datasets, respectively. We display the last 96 input steps based on each model’s optimal input length, along with the corresponding 96 predicted steps. Finally, LABEL:fig:app_vis_pems illustrates short-term forecasting for the PeMS03 dataset, where each model predicts 12 steps from a 96-step input.
@@ Appendix G Visualization
Model TimeBridge PDF (2024) Confidence
@@
ETTm1 $0.353\pm 0.014$ $0.388\pm 0.010$ $0.339\pm 0.008$ $0.372\pm 0.006$ $99\%$
ETTm2 $0.246\pm 0.004$ $0.310\pm 0.012$ $0.252\pm 0.003$ $0.313\pm 0.003$ $99\%$
ETTh1 $0.399\pm 0.010$ $0.420\pm 0.008$ $0.399\pm 0.015$ $0.419\pm 0.006$ $99\%$
ETTh2 $0.346\pm 0.018$ $0.394\pm 0.015$ $0.327\pm 0.009$ $0.376\pm 0.010$ $99\%$
Weather $0.218\pm 0.006$ $0.259\pm 0.004$ $0.225\pm 0.009$ $0.261\pm 0.006$ $99\%$
Electricity $0.149\pm 0.011$ $0.246\pm 0.007$ $0.159\pm 0.010$ $0.250\pm 0.015$ $99\%$
Traffic $0.360\pm 0.008$ $0.255\pm 0.013$ $0.383\pm 0.016$ $0.254\pm 0.010$ $99\%$
Solar $0.181\pm 0.002$ $0.239\pm 0.003$ $0.205\pm 0.008$ $0.265\pm 0.005$ $99\%$
Table 17: Standard deviation and statistical tests for TimeBridge and second-best method (PDF) on ETT, Weather, Electricity, Traffic, and Solar datasets.
Model TimeBridge TimeMixer (2024) Confidence

### v3

@@ E.2 Ablation Studies
We present the full results of the ablation studies discussed in the main text. Tab. 11 provides the complete results of the ablation on removing non-stationarity in both Integrated and Cointegrated Attention. Tab. 12 reports the full results on the impact and order of Integrated and Cointegrated Attention, with an illustrative visualization in Figure 6. Additionally, Tab. 13 shows the results of ablation on different modeling approaches for these attention mechanisms. Finally, Tab. 14 presents the results of varying the number of downsampled patches $M$ and its effect on forecasting performance.
@@
$\checkmark$ $\times$ 96 0.144 0.184 0.161 0.224 0.120 0.214 0.340 0.240
192 0.186 0.225 0.177 0.237 0.142 0.237 0.343 0.250
$\checkmark$ $\times$ 336 0.237 0.267 0.188 0.244 0.156 0.252 0.363 0.257
720 0.307 0.320 0.197 0.252 0.179 0.278 0.393 0.271
Avg. 0.219 0.249 0.181 0.239 0.149 0.245 0.360 0.255
@@ Appendix F Statistical Analysis
Model TimeBridge DeformableTST (2024) Confidence
@@
ETTm1 $0.344\pm 0.014$ $0.379\pm 0.010$ $0.348\pm 0.008$ $0.383\pm 0.006$ $99\%$
ETTm2 $0.246\pm 0.004$ $0.310\pm 0.012$ $0.257\pm 0.003$ $0.319\pm 0.003$ $99\%$
ETTh1 $0.397\pm 0.010$ $0.424\pm 0.008$ $0.404\pm 0.015$ $0.423\pm 0.006$ $99\%$
ETTh2 $0.341\pm 0.018$ $0.382\pm 0.015$ $0.328\pm 0.009$ $0.377\pm 0.010$ $99\%$
Weather $0.219\pm 0.006$ $0.249\pm 0.004$ $0.222\pm 0.009$ $0.262\pm 0.006$ $99\%$
Electricity $0.149\pm 0.011$ $0.245\pm 0.007$ $0.161\pm 0.010$ $0.261\pm 0.015$ $99\%$
Traffic $0.360\pm 0.008$ $0.255\pm 0.013$ $0.391\pm 0.016$ $0.278\pm 0.010$ $99\%$
Solar $0.181\pm 0.002$ $0.239\pm 0.003$ $0.185\pm 0.008$ $0.254\pm 0.005$ $99\%$
Climate $1.057\pm 0.012$ $0.494\pm 0.004$ $1.060\pm 0.011$ $0.496\pm 0.006$ $99\%$
Table 15: Standard deviation and statistical tests for TimeBridge and second-best method (DeformableTST) on ETT, Weather, Electricity, Traffic, Solar and Climate datasets.
Model TimeBridge TimeMixer (2024b) Confidence
@@
Appendix G Visualization
Figure 7 visualizes short-term fluctuations and long-term cointegration across stock sectors. Figure 8 provides additional examples of intra-variate attention maps comparing stationary and non-stationary conditions for different patches in the Electricity dataset. Figure 9 shows further examples of inter-variate attention maps in the Solar dataset under both stationary and non-stationary conditions. Figure 10, Figure 11, Figure 12, and Figure 13 present long-term forecasting visualizations for Weather, Solar, Electricity, and Traffic datasets, respectively. We display the last 96 input steps based on each model’s optimal input length, along with the corresponding 96 predicted steps. Finally, LABEL:fig:app_vis_pems illustrates short-term forecasting for the PeMS03 dataset, where each model predicts 12 steps from a 96-step input.
@@ Appendix G Visualization
Figure 13: Visualization of predictions from different models on the Traffic dataset.